OpenStudy (anonymous):

HELP PLSSSSSS what are the zeros of f(x)=x^2+5x+6

OpenStudy (jessicajohnson):

to find the zeros set the equation equal to zero first. 0= x^2 +5x+6 I like to rearrange so x is on the left hand side of equation x^2 +5x+6=0 then factor the equation (x+3)(x+2)=0 Now set each factor equal to zero x+3 =0 ; x+2=0 solve x=-3 ; x = -2